OpenStudy (anonymous):

perform the indicated operations then simplify each radical.... 26 sqrt 3a+5 sqrt 3 a - 9 sqrt 3a help!!!

OpenStudy (anonymous):

\[26\sqrt{3a} + 5 \sqrt{3a-9} \sqrt{3a}\]

OpenStudy (anonymous):